OpenStudy (anonymous):

Point A, located at (-8, 3), is the location of the storage closet and point C, located at (12, -4), is the director's office. Find point B, the location of the art supplies, if it is 2 over 5 the distance from point A to point C.

OpenStudy (anonymous):

@wolf1728 can you help me with this real quick

OpenStudy (anonymous):

@Compassionate can you help me please

OpenStudy (anonymous):

@whpalmer4 can you help me with this

OpenStudy (whpalmer4):

Point A is (-8,3) Point C is (12,-4) Find point B which is 2/5 of the way from A to C. Okay, this is easy. The x and y coordinates operate in a linear fashion, so the point 2/5 of the way from A to C is gotten by subtracting the x coordinate of A from the x coordinate of C (to get the distance between them), multiplying by 2/5, and adding back to the x coordinate of A. You do the same procedure for the y coordinates.

OpenStudy (anonymous):

ok thank you very much

OpenStudy (whpalmer4):

I'll check your answer when you're ready...

OpenStudy (anonymous):

is it -8,2.8

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

thank you for your help @whpalmer4

OpenStudy (whpalmer4):

sorry, had some connectivity issues. No, (-8,2.8) is not correct. Here's a picture

OpenStudy (anonymous):

then what is correct cause i did exactly what you said to do and got -8 and 2.8

OpenStudy (whpalmer4):

what is the x coordinate of A? what is the x coordinate of C?

OpenStudy (anonymous):

A is -8 C is 12

OpenStudy (whpalmer4):

Okay. What is the x coordinate of A subtracted from the x coordinate of C?

OpenStudy (anonymous):

-20

OpenStudy (whpalmer4):

No. 12 - (-8) =

OpenStudy (anonymous):

20

OpenStudy (whpalmer4):

Right. What's 2/5 of that?

OpenStudy (anonymous):

8

OpenStudy (whpalmer4):

and 8 plus the x coordinate of A =

OpenStudy (anonymous):

0

OpenStudy (whpalmer4):

right. now do that same procedure with the y coordinates

OpenStudy (anonymous):

i did it and got -4

OpenStudy (whpalmer4):

if you look at the graph I provided, you should be able to see that we went 2/5 of the way from A (upper left) to C (bottom right) along the x axis

OpenStudy (whpalmer4):

what's the y coordinate of C? what's the y coordinate of A?

OpenStudy (anonymous):

c is -4 and a is 3

OpenStudy (whpalmer4):

and \[-4-3=\]

OpenStudy (anonymous):

-7

OpenStudy (whpalmer4):

-7 * 2/5 =

OpenStudy (anonymous):

-2.8

OpenStudy (whpalmer4):

well, I prefer -14/5 but no matter. what is 3 + -2.8?

OpenStudy (anonymous):

0.2

OpenStudy (whpalmer4):

so the coordinates for B are?

OpenStudy (anonymous):

0,.2

OpenStudy (whpalmer4):

And does that point appear on the line drawn on my graph? By the way, you really need to put a 0 before the decimal point...you're just asking for trouble if you don't. And coordinates are put in parentheses to indicate that they are coordinates, not just numbers.

OpenStudy (anonymous):

yes it does gtg bye